We believe that 85% of the population of all Business Statistics students consider statistics to be an exciting subject. Suppose we randomly and independently selected 40 students from the population. If the true percentage is really 85%, find the probability of observing 39 or more students who consider statistics to be an exciting subject. Round to six decimal places.

0.001502

0.012107

0.987893

0.010604
Answer #1

x = no. of students who consider statistics to be an exciting subject

n =40

p=0.85

1-p = 0.15

P(x>=39) = P(39) + P(40)

= 40C39 * (0.85^39) * (0.15^1) + 40C40 * (0.85^40) * (0.15^0)

= 40 * (0.85^39) * (0.15^1) + 1 * (0.85^40) * (0.15^0)

= 0.012107

ANSWER : option (B) 0.012107

(please upvote)